DogGirl27:

What are the solution(s) of the quadratic equation 98 – x2 = 0? x = Plus or minus 2 StartRoot 7 EndRoot x = Plus or minus 6 StartRoot 3 EndRoot x = Plus or minus 7 StartRoot 2 EndRoot no real solution

GoodFellow:

This is easier, you can give it a try.

DogGirl27:

@goodfellow wrote:
This is easier, you can give it a try.
I cant do math very well, im doing online school and im struggling. If i knew what i was doing i wouldnt ask for help

Alex1513:

x = Plus or minus 7 StartRoot 2 EndRoot

Alex1513:

C is the answer

Extrinix:

Update of the answer choices to be viewed easier: \(A\) - \(x = \underline{+} 2\sqrt{7}\) \(B\) - \(x = \underline{+} 6\sqrt{3}\) \(C\) - \(x = \underline{+} 7\sqrt{2}\) \(D\) - \(no~real~solution\)

DogGirl27:

@alex1513 wrote:
C is the answer
thank you sm

GoodFellow:

I am sorry, I understand. Anyhow you should try to learn. You need to get an idea about the equation for solving quadratic equations.

Extrinix:

@princemagic wrote:
3
@alex1513 wrote:
C is the answer
suggestion as well as a rule, dont give direct answers.

princemagic:

so 3. 3 is C That's what i said first

DogGirl27:

thank you guys

Extrinix:

Here, I'll explain it.

Alex1513:

this is how it looks x = ±7√2

Extrinix:

\(98 – x^2 = 0\) get x by itself \(– x^2 = -98\) divide both sides by 1 \(x^2 = 98\) square 98 \(x = \underline{+}\sqrt{98}\) figure out 98 \(x = \underline{+}7\sqrt{2}\) answer